I have to write an equation to solve this problem... Not really sure how to begin.
"the length of a rectangle is 3 cm greater than its width. The perimeter is 24 cm. What are the dimensions?"

Answer by checkley77(12844)

The formula you need is:
P=2L+2W
L=W+3
24=2(W+3)+2W
24=2W+6+2W
24=4W+6
4W=24-6
4W=18
W=18/4
W=4.5 ANS. FOR THE WIDTH.
L=4.5+3=7.5 ANS. FOR THE LENGTH.
PROOF:
24=2*7.5+2*4,5
24=15+9
24=24